OpenStudy (anonymous):

If f(x) = −2x + 1, find f(−6).

OpenStudy (anonymous):

pug x=-6 and solve

OpenStudy (anonymous):

f(x) is your function, where f(x) is your output and x is your input. When it asks for f(-6), basically it's saying you have to input -6 wherever there is an x. In this case, it would look like this: f(-6) = -2(-6) + 1 f(-6) = 12 + 1 f(-6) = 13